What is consumption function?Suppose the consumption function of a society is as follows: C=100+0.85Y, where C=Total consumption expenditure and Y=Disposable income.
Answer the following questions:
a.What is the autonomous consumption?
b.What is the total consumption when disposable income is 5000?
c.What is induced consumption when disposable income is 8000?
d.Find MPS .Prove that MPC+MPS=1
1
Expert's answer
2015-01-15T09:21:52-0500
Consumption function is a single mathematical function used to express
consumer spending and could be written as follows:
C = A + MPC * D
where
C = Consumer spending
A = Autonomous consumption, or the level of consumption that would
still exist even if income was $0
MPC = Marginal propensity to consume, which is the ratio of
consumption changes to income changes
D = Real disposable income
a. So, autonomous consumption = A = 100
b. C=100+0.85Y = 100+0.85 * 5000 = 4350
c. Induced consumption is household consumption that varies with
income = MPC * D = 0.85 * 8000 = 6800
d. MPC+MPS=1
MPS = 1 – MPC = 1 – 0.85 = 0.15
We can prove it:
If Yd = C + S, then any change in Yd must resolve a change in C and
some part a change in S.
So we can write:
ΔYd = ΔC + ΔS
If we divide both sides by ΔYd, and we get:
1 = MPC + MPS